OpenStudy (meilendurcer):

How many different arrangements can be made with the letters from the word TOPIC? 3,125 10 24 120

OpenStudy (anonymous):

since all the letters are distinct, it is \(5!=5\times 4\times 3\times 2\)

OpenStudy (anonymous):

hope it is clear why 5 choices for the first letter 4 for the second 3 for the third 2 for the fourth 1 for the fifth by the counting principle you multiply number of the choices together

OpenStudy (meilendurcer):

thanks

OpenStudy (anonymous):

yw